OpenStudy (anonymous):

2(4w+5)+2W=32 is W = 3 ?

ganeshie8 (ganeshie8):

plug it and check

ganeshie8 (ganeshie8):

2(4w+5)+2W=32 2(4*3+5) + 2*3 = 32 2(12+5) + 6 = 32 2(17) + 6 = 32 34 + 6 = 32 40 = 32 FALSE. so W is not 3

OpenStudy (anonymous):

fine so what is it then ?

ganeshie8 (ganeshie8):

we will have to solve it and see

ganeshie8 (ganeshie8):

2(4w+5)+2W=32

ganeshie8 (ganeshie8):

distribute 2 into the parenthesis

ganeshie8 (ganeshie8):

2(4w+5)+2W=32 2*4W + 2*5 + 2W = 32 8W + 10 + 2W = 32

ganeshie8 (ganeshie8):

combine like terms, and solve W

OpenStudy (anonymous):

2.2

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

thank !

ganeshie8 (ganeshie8):

np :)